# Large Language Models in Qualitative Research: Can We Do the Data Justice?

### v1

@@ Abstract.
Qualitative researchers use tools to collect, sort, and analyze their data. Should qualitative researchers use large language models (LLMs) as part of their practice? LLMs could augment qualitative research, but it is unclear if their use is appropriate, ethical, or aligned with qualitative researchers’ goals and values. We interviewed twenty qualitative researchers to investigate these tensions. Many participants see LLMs as promising interlocutors with attractive use cases across the stages of research, but wrestle with their performance and appropriateness. Participants surface concerns regarding the use of LLMs while protecting participant interests, and call attention to an urgent lack of norms and tooling to guide the ethical use of LLMs in research. Given the importance of qualitative methods to human-computer interaction, we use the tensions surfaced by our participants to outline guidelines for researchers considering using LLMs in qualitative research and design principles for LLM-assisted qualitative data analysis tools.
@@
journalyear: 2024copyright: rightsretainedconference: Extended Abstracts of the CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: Extended Abstracts of the CHI Conference on Human Factors in Computing Systems (CHI EA ’25), April 26-May 1, 2025, Yokohama, Japandoi: 10.1145/3613905.3636301isbn: 979-8-4007-0331-7/24/05ccs: Human-centered computing HCI design and evaluation methodsccs: Human-centered computing HCI design and evaluation methods
@@ 1. Introduction
The qualitative research process is traditionally grounded in deep engagement with participants and their data through an iterative sensemaking process. Today, Large Language Models (LLMs), particularly those accessed through chat-based interfaces like ChatGPT  (OpenAI, 2024), are popular tools in research-related tasks, including in the social sciences (Grossmann et al., 2023). While a growing number of dedicated tools for qualitative research use LLMs to streamline segments of the qualitative research process, like scaffolding thematic analysis or automating qualitative coding  (Gebreegziabher et al., 2023; Gao et al., 2024; Choksi et al., 2024), direct interactions with ChatGPT and similar interfaces may be increasingly used by qualitative researchers as well, a topic which has not received thorough investigation to this point.
@@
LLMs distinguish themselves from prior tools for qualitative research for several reasons. First, LLMs are not dedicated qualitative analysis tools. Unlike most dedicated software platforms, users interact with them in natural language, leading to highly fluent interactions. LLMs go beyond simple document-organization or analysis functions in classic tools, and appear to facilitate highly flexible interpretations of documents, such as the ability to produce summaries or annotations of documents, behaviors which mimic aspects of the human sensemaking process core to qualitative research. Second, LLM use is currently dominated by one product, ChatGPT, that has rapidly entered general use by the public (Porter, [n. d.]), including qualitative researchers. Third, the dominant user interface metaphor (Neale and Carroll, 1997) of ChatGPT and its competitors is that of a jack-of-all-trades assistant and/or collaborator. This pattern is simultaneously intuitive thanks to the conversational interface, yet completely unguided. With wide-ranging and open-ended interactions with LLMs increasingly permeating everyday life, qualitative researchers are left to wonder whether or how these tools might apply to their research process or outputs. While some qualitative researchers support integrating LLMs into aspects of qualitative research, others, as  Soden et al. (2024) highlight, will be “concerned that increasingly all qualitative research, including that which draws on interpretive traditions of research design, is being evaluated from the perspective of positivism.”
@@
There is a critical and immediate need to understand the degree to which LLMs are shaping qualitative research processes. On the one hand, many researchers may benefit from the inclusion of LLM-based features in qualitative data analysis (QDA) systems, as has been explored in HCI literature (Gao et al., 2024; Jiang et al., 2021; Overney et al., 2024). However, the recent explosion of work using LLMs to annotate text in ways previously difficult for advanced models (Ziems et al., 2024; Bail, 2024) may risk the excitement of communities of scholars and tool builders without full consideration for potential tradeoffs to the sensemaking process. In particular, readily available tools that seek to automate the qualitative sensemaking process could decrease the amount of time researchers spend with their participant data, a potential risk to qualitative researchers’ depth of understanding.
@@
Furthermore, the speed of LLM development has outpaced guidance on their ethical use (Baronchelli, 2024). Most available policies focus on the danger of uploading sensitive information to cloud services (Ollion et al., 2024) or text generation in publications (Hosseini and Horbach, 2023). The HCI community specifically is contending with the need for developed policies that tackle the challenge of how to use AI ethically in research (Shen et al., 2023; Aubin Le Quéré et al., 2024). While guidelines are beginning to emerge, there are no guidelines specific to qualitative work today.
@@
This work critically reflects on the role of LLMs in qualitative research and offers guidance to researchers considering using LLMs. Through our work, we answer the following research questions:
@@
  • In what ways are LLMs being adopted into qualitative research processes today, and how might they be incorporated in the future?
@@
  • What do qualitative researchers say they stand to lose or gain from adopting LLMs into their research processes, and what tensions might arise?
@@
  • What are potential models for the ethical incorporation of LLMs into the qualitative research process?
@@
Qualitative research is a broad and rich tradition with scholars from numerous disciplines and ways of thinking. This paper focuses on the subset of researchers, including mixed-method researchers, who are considering using LLMs in their qualitative research processes. Those researchers should be clear-eyed and cautious about what is gained and lost when LLMs are used in data collection, analysis, or writing.
@@
This work surfaces both the curiosities and concerns of qualitative and mixed-method researchers across disciplines who are wrestling with what the introduction of LLMs means for their qualitative research process. Through twenty semi-structured interviews with researchers across academic communities, we sought to understand any current use, motivations for use, perceptions of usefulness, and concerns about the use of LLMs in qualitative research. Our findings show that qualitative researchers can envision leveraging LLMs for their work to help with a range of uses including generating interview materials and analyzing research artifacts. Specifically, the interactivity of ChatGPT has spurred researchers to think creatively about what chat-based interfaces can bring to their work. Simultaneously, researchers harbor fundamental concerns about ethics, unequal adoption of new technologies, model bias, and performance, leading some to abstain from the use of LLMs to conduct qualitative research and others to experiment despite concerns to participant privacy. New tools and practices are rapidly being adopted, but scholarly communities are not equipped for conducting ethical and effective research with them. As such, we are at a critical junction that requires urgent evaluation and recalibration of both our methodological approaches and thinking on research practices..
As a step forward, we outline guidelines for future qualitative researchers interested in using LLMs, and we present design opportunities and principles for LLM-based qualitative data analysis tools.
@@ 2. Related Work
Qualitative research is a broad term with different definitions and applications across disciplines (Hammersley, 2013). In the late 1960s, the method first emerged in retaliation to quantitativeand positivist approaches to science characterized by hypothesis-testing and objectivity (Hammersley and Richardson, 1996). HCI scholars often turn to qualitative methods when studying populations or contexts that require an in-depth exploration of experiences, behaviors, and interactions (Randazzo and Ammari, 2023; Randazzo et al., 2023; Rankin and Han, 2019; Freeman and Wohn, 2020; Bardzell and Bardzell, 2011b, a).
@@
As a field, HCI hosts strong qualitative approaches but also mixed-method research, which has sometimes led to tensions in how contributions should be evaluated in the field. In particular, a stigma surrounding the perceived rigor of qualitative research (Charmaz and Thornberg, 2021) has sparked debates over how best to evaluate such methods (Kvale, 1995; McDonald et al., 2019; Rolfe, 2006). As a result, cting qualitative research can involve a complex and sometimes creative negotiation with the norms and evaluative criteria of one’s research communities (Denzin et al., 2024). Within the HCI community, there have been concerns that increasingly quantitative evaluation techniques, such as providing participant demographics, are being unduly required of qualitative scholars (Soden et al., 2024). LLMs may lend the impression that qualitative inquiry can be automated, and their integration into Qualitative Data Analysis (QDA) software may increasingly impose positivist approaches that conflict with interpretive traditions.
@@
Approaches to analyzing qualitative data (e.g., grounded theory, thematic analysis) were originally designed for manual, pen-and-paper applications (Deterding and Waters, 2021). However, the widespread adoption of software tools for qualitative data analysis marked a significant and, at times, contentious shift in the field (Gilbert et al., 2014). Early critiques of qualitative software centered around concerns that it created too much distance from the data, compromised the depth of analysis (Blismas and Dainty, 2003), or could be misused by researchers who lacked a reflective understanding of them (Gilbert et al., 2014).
@@
Advancements in natural language processing and machine learning over the last decades have widened possibilities for data analysis. LLMs are quickly being adapted for text annotation in increasingly interpretive contexts (Ziems et al., 2024; Bail, 2024). In HCI, tools that leverage these advancements have been developed to assist qualitative researchers since before the most recent LLM era. Past published work has surfaced qualitative HCI researchers’ imagined uses of AI in analysis, and found researchers were open to using AI in data exploration and annotation with stipulations regarding researcher agency (Feuston and Brubaker, 2021) or to “support serendipity” in the analysis process (Jiang et al., 2021), while also surfacing concerns about potential tradeoffs for qualitative research these tools could present. Past work has also considered the tensions and potential compatibilities between computational tools and interpretive work (Baumer et al., 2017; Nelson, 2020). More recent tools have built on this work, leveraging a new generation of LLMs to create tools for qualitative researchers that bear researcher goals in mind, like agency over the coding process (Overney et al., 2024) or convergent discussion on coding among collaborators (Gao et al., 2024). Each contains insights from users that give insight into the promises and pitfalls of LLMs in qualitative research, but evaluations focus on users’ perceptions of particular tools. Given the sudden, widespread ubiquity of LLMs through task-agnostic interfaces like ChatGPT, experiences with AI tools have become more common for all kinds of users, including those not in computing. The sudden uptake of LLMs in society has also changed common perceptions about which uses cases may be possible in the future, motivating a fresh examination of current uses and attitudes towards LLMs in qualitative research that also includes researchers outside of HCI.
@@ 3. Methods
To gain a deeper understanding of the experiences qualitative researchers are having with generative AI tools, we conducted a series of twenty semi-structured interviews in June-August 2024. We analyzed the interviews using inductive thematic analysis, inspired by grounded theory  (Braun and Clarke, 2006), in which concepts emerge from the data (Jaccard and Jacoby, 2020). These interview procedures were approved by the Institutional Review Board at [anonymous institutions].
@@
Participants were recruited mostly based on personal networks, past work, and through in-person events. Members of the research team also participated in in-person events, where they were able to conduct interviews on-site. On some occasions, we encountered participants’ prior works online and emailed them to request an interview. The inclusion criterion for this interview was simple: we sought to speak with researchers who consider themselves qualitative scholars, including researchers who use qualitative methods as part of a mixed methods approach to their work. Since we sought to maximize variety of perspectives, we did not recruit participants based on their use or knowledge of LLMs, and included participants across academic domains. Since qualitative research perspectives are field-dependent, we considered it important to allow potential interviewees to self-define as qualitative researchers. These recruitment strategies resulted in twenty participant interviews, summarized in Table 1.
The interview protocol was designed to elicit perspectives on LLMs in qualitative research, with a focus on data collection and analysis processes, as they are more unique to qualitative research than other steps of the research process, which have been the focus of other work (Kapania et al., 2024). For this study, data collection was defined as any part of the research process that contributes towards how researchers obtain data from participants. LLMs may play an active role in data collection, e.g. through creating recruitment materials for different formats. Data analysis was defined as any part of the research process that relates to the analysis of gathered data. LLMs may play a role in data analysis, e.g. through applying qualitative codes. We opened the interview with questions about the researcher’s qualitative work, practices, and their general familiarity with large language models. We then asked questions about data collection and analysis, asking the researcher if they have or could imagine using LLMs for any portion of these research processes. If they said they had used LLMs for that purpose, we asked them about their motivations, experiences, and ethical considerations. If they had not used LLMs before for that purpose, we asked them why, and whether they had major concerns with others using LLMs for these purposes. We concluded our interviews by discussing the participants’ perceptions of LLMs in qualitative research, soliciting potential use cases and concerns for using LLMs in qualitative research. The full protocol is included in the Appendix.
We analyzed the data using inductive thematic analysis techniques inspired by grounded theory (Braun and Clarke, 2006). Once the interviews were recorded, three researchers open coded two interviews each, and each compiled a proposed codebook structure, which we consolidated through discussion. Two further interviews were fully coded with this first-round codebook, and we iterated once more on the codebook to yield the final codebook. We then divided the twenty interviews among coders, and met regularly to discuss the process. We used Dedoose for collaborative code application to interview transcripts. We primarily avoided using LLMs for this research, with the exception of finding specific works of related literature, which we then verified externally. To protect the identity of participants, any specific research studies or topics have been changed.
We take a moment to reflect on how our own positionality influenced this study. The lead authors of this study are mixed-method researchers in HCI and communication. In our own work, we have felt a desire and some pressure to incorporate LLMs into our research processes, while being unsure how to ethically navigate adopting this technology. This work is an effort to tackle these difficulties head-on and to shine a light on the value of qualitative work in HCI.

### v2

# Large Language Models in Qualitative Research: Uses, Tensions, and Intentions

@@ Abstract.
Qualitative researchers use tools to collect, sort, and analyze their data. Should qualitative researchers use large language models (LLMs) as part of their practice? LLMs could augment qualitative research, but it is unclear if their use is appropriate, ethical, or aligned with qualitative researchers’ goals and values. We interviewed twenty qualitative researchers to investigate these tensions. Many participants see LLMs as promising interlocutors with attractive use cases across the stages of research, but wrestle with their performance and appropriateness. Participants surface concerns regarding the use of LLMs while protecting participant interests, and call attention to an urgent lack of norms and tooling to guide the ethical use of LLMs in research. We document the rapid and broad adoption of LLMs across surfaces, which can interfere with intentional use vital to qualitative research. We use the tensions surfaced by our participants to outline recommendations for researchers considering using LLMs in qualitative research and design principles for LLM-assisted qualitative research tools.
@@
copyright: acmcopyrightjournalyear: 2025conference: CHI Conference on Human Factors in Computing Systems; April 26-May 1, 2025; Yokohama, Japanbooktitle: CHI Conference on Human Factors in Computing Systems (CHI ’25), April 26-May 1, 2025, Yokohama, Japandoi: 10.1145/3706598.3713726isbn: 979-8-4007-1394-1/25/04ccs: Human-centered computing HCI design and evaluation methodsccs: Human-centered computing Empirical studies in HCI
@@ 1. Introduction
The qualitative research process is grounded in deep engagement with participants and their data through an iterative sensemaking process. Today, Large Language Models (LLMs), particularly those accessed through chat-based interfaces like ChatGPT  (OpenAI, 2024), are popular tools in research-related tasks, including within the social sciences (Grossmann et al., 2023; Liao et al., 2024). While a growing number of dedicated tools for qualitative work use LLMs to streamline research processes (Gebreegziabher et al., 2023; Gao et al., 2024; Choksi et al., 2024), direct interactions with ChatGPT and similar interfaces may also be increasingly used by qualitative researchers, a topic which has not received thorough investigation to this point.
@@
LLMs distinguish themselves from prior tools for qualitative research. First, chat-based LLMs are marketed as general purpose tools, not dedicated qualitative analysis tools. Users interact with them in natural language, unlike most dedicated software platforms. LLMs go beyond the simple document-organization or analysis functions in classic tools, and facilitate highly flexible interpretations of documents, such as the ability to produce summaries or annotations, behaviors which mimic aspects of the human sensemaking process. Second, LLM use is currently dominated by one product, ChatGPT, that has rapidly entered general use by the public (Porter, 2023), including by qualitative researchers. Third, the dominant user interface metaphor (Neale and Carroll, 1997) of ChatGPT and its competitors is that of a jack-of-all-trades assistant and/or collaborator. This pattern is intuitive thanks to the conversational interface, yet completely unguided. With wide-ranging and open-ended interactions with LLMs increasingly permeating everyday life, qualitative researchers are left to wonder whether or how these tools might apply to their research process or outputs. While some qualitative researchers support integrating LLMs into aspects of qualitative research, others, as  Soden et al. (2024) highlight, will be “concerned that increasingly all qualitative research, including that which draws on interpretive traditions of research design, is being evaluated from the perspective of positivism.”
@@
There is a critical and immediate need to understand the degree to which LLMs are shaping qualitative research processes. On the one hand, many researchers may benefit from the inclusion of LLM-based features in qualitative data analysis (QDA) systems, as has been explored in HCI literature (Gao et al., 2024; Jiang et al., 2021; Overney et al., 2024). However, the recent explosion of work using LLMs to annotate text in more complex ways than previously possible (Ziems et al., 2024; Bail, 2024) may risk the excitement of communities of scholars and tool builders without full consideration for potential tradeoffs to the sensemaking process. In particular, readily available tools that seek to automate the qualitative sensemaking process could decrease the amount of time researchers spend with their participant data, a potential risk to qualitative researchers’ depth of understanding.
@@
Furthermore, the speed of LLM development has outpaced guidance on their ethical use (Baronchelli, 2024). Most policies currently guiding use focus on the danger of uploading sensitive information to cloud services (Ollion et al., 2024) or text generation in publications (Hosseini and Horbach, 2023). The HCI community is contending with the need for developed policies that tackle how to use AI ethically in research (Shen et al., 2023; Aubin Le Quéré et al., 2024). While guidelines are beginning to emerge, there are no guidelines specific to qualitative work today.
@@
This work critically reflects on the role of LLMs in qualitative research and offers considerations and recommendations for researchers considering using LLMs. Through our work, we answer the following research questions:
@@
  • RQ1: In what ways are LLMs being adopted into qualitative research processes today, and how might they be incorporated in the future?
@@
  • RQ2: What do qualitative researchers say they stand to lose or gain from adopting LLMs into their research processes, and what tensions might arise?
@@
  • RQ3: What are potential models for the ethical incorporation of LLMs into the qualitative research process?
@@
Qualitative research is a broad and rich tradition with scholars from varied disciplines and ways of thinking. This paper focuses on the subset of researchers, including mixed-method researchers, who are considering using LLMs in their qualitative research processes. Those researchers should be clear-eyed about what is gained and lost when LLMs are used in data collection, analysis, or writing.
@@
This work surfaces both the curiosities and concerns of qualitative and mixed-method researchers across disciplines who are wrestling with what the introduction of LLMs means for their qualitative research process. Through twenty semi-structured interviews with researchers across academic communities, we sought to understand any current use, motivations for use, perceptions of usefulness, and concerns about the use of LLMs in qualitative research. Our findings show that qualitative researchers can envision leveraging LLMs for their work to help with a range of uses including generating interview materials and analyzing research artifacts. Specifically, the interactivity of ChatGPT has spurred researchers to think creatively about what chat-based interfaces can bring to their work. Simultaneously, researchers harbor fundamental concerns about ethics, unequal adoption of new technologies, model bias, and performance, leading some to abstain from the use of LLMs to conduct qualitative research, and others to experiment despite concerns regarding participant privacy. New tools and practices are rapidly being adopted, but scholarly communities are not yet equipped to conduct ethical and effective research with them. As such, we are at a junction that urgently requires thorough evaluation and recalibration of both our methodological approaches and our thinking on research practices. As a step forward, we outline considerations and recommendations for future qualitative researchers interested in using LLMs, and we present design opportunities and principles for creating LLM-based qualitative data analysis tools given our findings.
@@ 2. Related Work
Qualitative research is a broad term with different definitions and applications across disciplines (Hammersley, 2013). In the late 1960s, the method emerged in retaliation to quantitative and positivist approaches to science characterized by hypothesis testing and objectivity (Hammersley and Richardson, 1996). HCI scholars often turn to qualitative methods when studying populations or contexts that require deep exploration and understanding of experiences, behaviors, and interactions (Randazzo and Ammari, 2023; Randazzo et al., 2023; Rankin and Han, 2019; Freeman and Wohn, 2020; Bardzell and Bardzell, 2011b, a).
@@
As a field, HCI is both home to researchers with qualitative approaches as well as mixed-method researchers, a fact which has sometimes led to tension in the community regarding how contributions should be evaluated. In particular, a stigma surrounding the perceived rigor of qualitative research (Charmaz and Thornberg, 2021) has sparked debates over how best to evaluate such methods (Kvale, 1995; McDonald et al., 2019; Rolfe, 2006). As a result, conducting qualitative research can involve a complex and sometimes creative negotiation with the norms and evaluative criteria of one’s research communities (Denzin et al., 2024). Within the HCI community, there have been concerns that increasingly quantitative evaluation techniques, such as providing participant demographics, are being unduly required of qualitative scholars (Soden et al., 2024). LLMs may lend the impression that qualitative inquiry can be automated, and their integration into Qualitative Data Analysis (QDA) software may increasingly impose positivist approaches that conflict with interpretivist traditions.
@@
Approaches to analyzing qualitative data (e.g., grounded theory, thematic analysis) were originally designed for manual, pen-and-paper applications (Deterding and Waters, 2021). In the long history of qualitative research practice, many new technologies have been introduced, including the recording and digitization of interview data, popularization of automated transcription, and introduction of phone interviews, eventually followed by online interviews (Brown et al., 2002; Bryda and Costa, 2023). The widespread adoption of software tools for qualitative data analysis marked a significant and, at times, contentious shift in the field (Gilbert et al., 2014). Early critiques of qualitative software centered around concerns that these tools created too much distance from the data (Gilbert, 2002), compromised the depth of analysis (Blismas and Dainty, 2003), or could be misused by researchers who lacked an understanding of them (Gilbert et al., 2014).
@@
Advancements in natural language processing (NLP) and machine learning have widened methods for data analysis over the past several decades, and the sudden rise of LLMs has, once again, opened up new possibilities. LLMs are being used for text annotation in interpretive contexts (Ziems et al., 2024; Bail, 2024), including qualitative coding (Choksi et al., 2024; Dunivin, 2024). Past work has considered both the potential tensions and compatibilities between computational tools and interpretive work, grounding the conversation regarding the potential usefulness of computational tools in ways that can support qualitative work (Baumer et al., 2017; Nelson, 2020). In HCI, tools that leverage advancements in machine learning have been developed to assist qualitative researchers with data and text analysis since before the recent LLM era. The new generation of LLMs has ushered in new kinds of tools, with developers creating tools for qualitative researchers that bear researcher goals in mind, like agency over the coding process (Overney et al., 2024), convergent discussion on coding among collaborators (Gao et al., 2024), or incorporating trauma-informed design principles (Tseng et al., 2025). Studies introducing and evaluating these tools contain insights regarding the promises and pitfalls of LLMs in qualitative research, but largely focus on users’ perceptions of the particular tools being developed for a task-specific use case, such as data annotation. Given the sudden, widespread ubiquity of LLMs through task-agnostic interfaces like ChatGPT, experiences with AI tools have become more common while being less scaffolded by purpose-built software, warranting investigation of current LLM use by qualitative researchers in a task-agnostic context.
Beyond the evaluation of specific software interfaces, prior work has also explored the concept of using AI tools to augment qualitative work. Jiang et al. (2021) and Feuston and Brubaker (2021) tackle the question of how qualitative researchers may respond to incorporating AI in their workflows. Jiang et al. (2021) identify that AI tools for qualitative work should “honor serendipity, human agency, and ambiguity.” Feuston and Brubaker (2021) highlight that intentional use and ordering of AI tools in qualitative tasks is paramount, and that the use of any computational tools may entail shifts in qualitative processes towards scale, abstraction, and delegation. These works lay key groundwork for this study, but were both published prior to the popular adoption of LLMs. Both papers presuppose the use of AI in the qualitative research workflow in bounded, task-specific ways, an idea which largely contrasts with the fluid ways that researchers often use LLMs throughout workflows today. The sudden adoption of LLMs warrants an urgent re-examination of AI tools in qualitative work, with early research suggesting that 81% of researchers use LLMs in their research practices (Liao et al., 2024). Many papers have examined the application of an LLM, or of ChatGPT in particular, to the problem of a thematic analysis for an applied case, with examples across fields as varied as medicine and education (Pentucci et al., 2024). Researchers often report promise regarding LLMs’ ability to capture some or most themes humans do, but also perceive a lack of depth, nuance, or variety in identified themes (Spangler et al., 2024; Morgan, 2023; Jalali and Akhavan, 2024; Li et al., 2024; Pentucci et al., 2024; Wachinger et al., 2024). Simultaneously, some work raises or notes researchers’ concerns regarding the reliability and validity of using LLMs in qualitative research, and privacy concerns that may arise when using them (Yan et al., 2024; Friedman et al., 2024). As LLMs permeate society, sociotechnical gaps (Ackerman, 2000) may arise from the use of popular tools like ChatGPT for qualitative research purposes, where the social expectations of research do not match the underlying technical capabilities (Liao and Xiao, 2023; Ehsan et al., 2023). In light of these societal shifts, we endeavor a fresh examination of the current uses, attitudes, and tensions that arise when using LLMs in qualitative research. Through highlighting a range of considerations for qualitative researchers to scaffold their intentional use of these tools, we chart a path forward for using LLMs conscientiously in qualitative work.
@@ 3. Methods
To understand how qualitative researchers are interacting with generative AI tools or choosing not to, we conducted a series of twenty semi-structured interviews in June-August 2024. We analyzed the interviews using inductive thematic analysis, inspired by grounded theory  (Braun and Clarke, 2006), in which concepts emerge from the data (Jaccard and Jacoby, 2020). These interview procedures were approved by the Institutional Review Board at Cornell University.
@@
Participants were recruited through personal networks, email outreach about past work, and through in-person events and conferences. The inclusion criterion for this interview was simple: we sought to speak with researchers who consider themselves qualitative scholars, including mixed-method researchers. Since we wanted to include a variety of perspectives, we did not recruit participants based on their use or knowledge of LLMs, and we recruited participants across academic domains. Since qualitative research perspectives can vary by field, we considered it important to allow potential interviewees to self-define as qualitative researchers.
These recruitment strategies resulted in twenty participant interviews, summarized in Table 1. Our participants spanned seven academic fields, include sixteen PhD students, and are balanced between nine primarily qualitative and eleven mixed-method scholars. Beyond this high-level information, we did not explicitly collect demographic information from participants. Nonetheless, we note that a majority of participants presented as women and as researchers of color, and five participants referenced being non-native English speakers in interviews. Outreach in our personal networks resulted in a population that was largely located in North American Universities, with one exception.
@@
The interview protocol was designed to elicit perspectives on LLMs in qualitative research, with a focus on the data collection and analysis processes. For this study, data collection was defined as any part of the research process that contributes towards how researchers obtain data from participants. Data analysis was defined as any part of the research process that relates to the analysis of gathered data. We opened the interview with questions about the researcher’s qualitative work, practices, and their general familiarity with large language models. We then asked questions about data collection and analysis, asking the researcher if they have or could imagine using LLMs for any portion of these research processes. If they said they had used LLMs for that purpose, we asked them about their motivations for doing so, experiences of doing so, including perceptions of model performance, and ethical considerations during the process. If they had not used LLMs for the steps we outline, we asked them why they had not, and whether they had concerns with others using LLMs for these purposes. We concluded by discussing the participants’ overall perceptions of LLMs in qualitative research, and soliciting both their perceived potential use cases as well as their concerns. Interviews lasted 42 minutes on average (min: 24, max: 59). The full interview protocol is included in the Appendix.
We iteratively analyzed the data using inductive thematic analysis techniques inspired by grounded theory (Braun and Clarke, 2006). While data collection was ongoing, three researchers open coded two interviews each, and each compiled a proposed codebook centered around identifying main themes, which were consolidated into one hierarchical codebook. These initial themes motivated us to balance recruitment between computing and non-computing scholars, and participants with varying approaches to qualitative research. Data collection continued until we obtained a balance of academic domains and research approaches, and the main identified themes remained stable indicating theoretical saturation (Glaser and Strauss, 2017). Using our first-round codebook as a starting point, two further interviews were fully coded by three researchers each, and the codes were refactored to yield the final codebook. We then divided the twenty interviews among coders, and met regularly to discuss the process and update codes where needed. The two primary authors re-reviewed each transcript and the applied codes for thoroughness, verifying with each other when uncertain. Throughout the analysis process, we also constructed various artifacts, including interview notes, memos, and tables that summarized key context about each participant. We used Dedoose for collaborative code application to interview transcripts. We primarily avoided using LLMs for this research, with the exception of finding specific works of related literature, which we then verified externally. To protect the identity of participants, any specific research studies or topics have been changed.
We reflect on how our own positionality influenced this study. The authors are mixed-method researchers in HCI, NLP, and communication. In our own work, we have felt interest in incorporating LLMs into our research processes, while being unsure how to ethically navigate adopting this technology. This work is an effort to consider these opportunities and challenges head-on and to shine a light on the value of qualitative work in HCI.